# Factor Graph Optimization of Error-Correcting Codes for Belief Propagation Decoding

### v1

@@ Abstract
The design of optimal linear block codes capable of being efficiently decoded is of major concern, especially for short block lengths. As near capacity-approaching codes, Low-Density Parity-Check (LDPC) codes possess several advantages over other families of codes, the most notable being its efficient decoding via Belief Propagation. While many LDPC code design methods exist, the development of efficient sparse codes that meet the constraints of modern short code lengths and accommodate new channel models remains a challenge. In this work, we propose for the first time a data-driven approach for the design of sparse codes. We develop locally optimal codes with respect to Belief Propagation decoding via the learning on the Factor graph (also called the Tanner graph) under channel noise simulations. This is performed via a novel tensor representation of the Belief Propagation algorithm, optimized over finite fields via backpropagation coupled with an efficient line-search method. The proposed approach is shown to outperform the decoding performance of existing popular codes by orders of magnitude and demonstrates the power of data-driven approaches for code design.
@@ 1 Introduction
Reliable digital communication is of major importance in the modern information age and involves the design of codes that can be robustly and efficiently decoded despite noisy transmission channels. Over the last half-century, significant research has been dedicated to the study of capacity-approaching Error Correcting Codes (ECC) [58]. Despite the initial focus on short and medium-length linear block codes [5], the development of long channel codes [22, 17] has emerged as a viable approach to approaching channel capacity [6, 42, 56, 57, 2, 38, 35].
While the NP-hard maximum likelihood rule defines the target decoding of a given code, developing more practical solutions generally relies on theories grounded upon asymptotic analysis over conventional communication channels. However, modern communication systems rely on the design of short and medium-block-length codes [37] and the latest communication settings provide new types of channels. This is mainly due to emergent applications in the modern wireless realm requiring the transmission of short data units, such as remote command links, Internet of Things, and messaging services [18, 7, 51, 19, 21]. These challenges call for the formulation of data-driven solutions, capable of adapting to various settings of interest and constraints, generally uncharted by existing theories.
The vast majority of existing machine-learning solutions to the ECC problem concentrate on the design of neural decoders. The first neural models focused on the implementation of parameterized versions of the legacy Belief Propagation (BP) decoder [44, 45, 40, 46, 9]. Recently, state-of-the-art learning-based de novo decoders have been introduced, borrowing from well-proven architectures from other domains. A Transformer-based decoder that incorporates the code into the architecture has been recently proposed by [12], outperforming existing methods by sizable margins and at a fraction of their time complexity. This architecture has been subsequently integrated into a denoising diffusion models paradigm, further improving results [11]. Subsequently, a universal neural decoder has been proposed in [14], capable of unified decoding of codes from different families, lengths, and rates. Most recently and related to our work, [15] developed an end-to-end learning framework capable of co-learning binary linear block codes along with the neural decoder.
@@
In this work, given the ubiquity and advantages of the Belief Propagation (BP) algorithm [52, 56] for sparse codes, we consider the optimization of codes with respect to BP via the learning of the underlying factor/Tanner graph. As far as we can ascertain, this is the first time a data-driven solution is given for the design of the codes themselves based on a classical decoder. Such a solution induces a very low overhead (if any) for integration in the existing decoding solutions.
@@ 2 Related Works
Neural decoder or data-driven contributions generally focus on short and moderate-length codes for two main reasons. First, classical decoders reach the capacity of the channel for large codes, and second, the emergence of applications driven by the Internet of Things requires effective decoders for short to moderate-length codes. For example, 5G Polar codes have code lengths of 32 to 1024 [37, 21].
@@
Previous work on neural decoders is generally divided into two main classes: model-free and model-based. Model-free decoders employ general types of neural network architectures [10, 24, 33, 4, 29, 12, 11, 14, 13]. Model-based decoders implement parameterized versions of classical Belief Propagation (BP) decoders, where the Tanner graph is unfolded into an NN in which scalar weights are assigned to each variable edge. This results in an improvement in comparison to the baseline BP method for short codes [44, 46, 53, 54, 36]. While model-based decoders benefit from a strong theoretical background, the architecture is overly restrictive, which generally enforces its coupling with high-complexity NN [47]. Also, the improvement gain generally vanishes for more iterations and longer codewords [26] and the integration cost remains very high due to both computational and memory requirements.
@@
While neural decoders show improved performance in various communication settings, there has been very limited success in the design of novel neural coding methods, which remain impracticable for deployment [50, 32, 30]. Recently, [15] provided a new differentiable way of designing binary linear block codes (i.e., parity-check matrices) for a given neural decoder also showing improved performance with classical decoders.
@@
Belief-propagation decoding has multiple advantages for LDPC codes [23, 57, 56]. A large number of LDPC code (parity check matrix) design techniques exist in the literature, depending on the design criterion. Among them, Gallager [23] developed the first regular LDPC codes as the concatenations of permuted sub-matrices. MacKay [43] demonstrated the ability of sparse codes to reach near-capacity limits via semi-randomly generated matrices. Irregular LDPC codes have been developed by [56, 38, 16] where the decoding threshold can be optimized via density-evolution. Progressive Edge Growth [27, 28] has been proposed to design large girth codes. Certain classes of LDPC array codes have been presented in [20] and LDPC codes with combinatorial design constraints have been developed in [60]. Finite geometry codes have been developed in [39, 34] and repeat-accumulate codes have been proposed by [62, 31, 48]. However, the classical methods are not data-driven and are difficult to adapt to the design of codes under constrained settings of interest (e.g., short codes, modern channels, structure constraints, etc.).
@@ 3 Background
We assume a standard transmission protocol using a linear block code $C$. The code is defined by a generator matrix $G\in\{0,1\}^{k\times n}$ and the parity check matrix $H\in\{0,1\}^{(n-k)\times n}$ is defined such that $GH^{T}=0$ over the order 2 Galois field $GF(2)$. The parity check matrix $H$ entails what is known as a Tanner graph [59], which consists of $n$ variable nodes and $(n-k)$ check nodes. The edges of this bipartite graph correspond to the on-bits in each column of the matrix $H$.
The input message $m\in\{0,1\}^{k}$ is encoded by $G$ to a codeword $c\in C\subset\{0,1\}^{n}$ satisfying $Hc=H(Gm)=0$ and transmitted via a Binary-Input Symmetric-Output channel, e.g., an AWGN channel. Let $y$ denote the channel output represented as $y=c_{s}+\varepsilon$, where $c_{s}$ denotes the transmission modulation of $c$ (e.g., Binary Phase Shift Keying (BPSK)), and $\varepsilon$ is random noise independent of the transmitted $c$. The main goal of the decoder $f_{H}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ conditioned on the code (i.e., $H$) is to provide a soft approximation $\hat{x}=f_{H}(y)$ of the codeword.
@@
Here, $v\in\{1,\dots,n\}$ denotes the index corresponding to the $v^{th}$ element of the channel output $y$, for the corresponding bit $c_{v}$ we wish to recover.
@@
Here, each variable node is indexed by the edge $e=(c,v)$ on the Tanner graph and $N(v)=\{(c,v)|H(c,v)=1\}$, i.e, the set of all edges in which $v$ participates. By definition $x^{0}=0$ such that the messages are directly determined by the vector $L$ for $k=0$.
@@
Figure 1: For the Hamming(7,4) Code: (a) parity check matrix, (b) Tanner graph, and (c) the corresponding Trellis graph with two iterations, with odd layers in blue and even layers in red.
@@ 4 Method
The performance of BP is strongly tied to the underlying Tanner graph induced by the code. BP and its variants are generally implemented over a fixed sparse graph, such that the only degree of freedom resides in the number of decoding iterations. While several recent contributions [44, 46] aim to enhance the BP algorithm by augmenting the Trellis graph with neural networks, these approaches assume and maintain fixed codes. Here, we propose optimizing the code for the BP algorithm on a decoding setting of interest. Given a trainable binary parity check matrix $H$, we wish to obtain BP-optimized codes by solving the following parameterized optimization problem
@@
Here, $G(H)$ denotes the generator matrix defined injectively by $H$, $\phi$ denotes the modulation function such that $c_{s}=\phi(c)$, and $\mathcal{Z}$ is the channel noise distribution. $f_{H,T}$ denotes the BP decoder built upon $H$ with $T$ iterations, $\mathcal{D}$ denotes the distance metric of interest, and $\mathcal{R}$ denotes the potential hard/soft regularization of interest, e.g., sparsity or constraints on the code structure.
@@
To obtain BP codes, we propose a Tanner graph learning approach, where the bipartite graph is assumed as complete with binary-weighted edges. This way, the tensor reformulation of BP weighted by $H$ allows a differentiable optimization of the Tanner graph itself. The two alternating stages of BP can now be represented in a differentiable matrix form rather than its static graph formulation, where the variable layers can be rewritten as
$Q_{ij}$ $=L+\sum_{j^{\prime}\in C_{i}\setminus j}{R_{j^{\prime}i}}\equiv L+\sum_{j^{\prime}}{R_{j^{\prime}i}H_{j^{\prime}i}}-R_{ji}\,,$ (5)
@@
$=2\text{arctanh}\left(\prod_{i^{\prime}}\left({\tanh\left(\frac{Q_{i^{\prime}j}H_{ji^{\prime}}}{2}\right)+\left(1-H_{ji^{\prime}}\right)}\right)/\left(\tanh\left(\frac{Q_{ij}}{2}\right)+\epsilon\right)\right)\,,$
where $C_{i}$ and $V_{j}$ correspond to the non-zero elements in column $i$ and row $j$ of $H$, respectively, while the ones elements in $(1-H)\in\{0,1\}^{(n-k)\times n}$ satisfy the identity element of multiplication. As we can observe, BP remains differentiable with respect to $H$ as a composition of differentiable functions.
@@ 7             R = 2*atanh( prod(tmp*H+(1-H),dim=1)/(tmp+eps) )
The tensor reformulation solves the major challenge of graph learning (challenge (iv)). Challenges (ii) and (iii) are also eliminated in our formulation. First, since for any given $H$ the conditional independence of error probability under symmetry [57] is satisfied for message passing algorithms, it is enough to optimize the zero codeword only, i.e., $c=Gm=0$, removing then any dependency on $G$ in the objective (challenge (ii)). As a byproduct, we obtain that the optimization problem is invariant to the choice of modulation, whether differentiable or not (challenge (iii)).
To optimize $H$ (challenge (i)) we relax the NP-hard binary programming problems to an unconstrained objective where, given a parameter matrix $\Omega\in\mathbb{R}^{(n-k)\times n}$, we have $H\coloneqq H(\Omega)=\operatorname{bin}(\Omega)$. Here $\operatorname{bin}(\cdot)$ refers to the element-wise binarization operator implemented via the shifted straight-through-estimator (STE) [3] defined such that
@@
where $c_{s}=\phi(c)$ denotes the modulated zero codeword and $\varepsilon_{i}$ denotes the $i^{th}$ noise sample drawn from the channel noise distribution.
While highly non-convex, the objective is (sub)differentiable and thus optimizable via classical first-order methods. Since $H$ is binary, only changes in the sign of $\Omega$ are relevant for the optimization, so most gradient descent iterations remain ineffective in reducing the objective using conventional small learning-rate regimes. Thus, given the gradient $\nabla_{\Omega}\mathcal{L}$ computed on sufficient statistics, we propose a line-search procedure capable of finding the optimal step size.
In general, efficient line search methods [49] assume local convexity or a smooth objective [61]. Since this is not our case, we propose a novel efficient grid-search approach optimized to our binary programming setting. While classical grid search methods look for the optimal step size on handcrafted predefined sample points, in our binary setting we can search only for the step sizes inducing a flip of the sign in $\Omega$, limiting the maximum number of relevant grid samples to $n(n-k)$. Thus, the line-search problem is now given by
$\lambda^{*}=\operatorname*{arg\,min}_{\lambda\in\mathcal{I}_{\Omega}}\mathcal{L}(\Omega-\lambda\nabla_{\Omega}\mathcal{L}),\ \ \ \ \ \ \ \mathcal{I}_{\Omega}=\{s_{i}=\bigg{(}\frac{\Omega}{\nabla_{\Omega}\mathcal{L}}\bigg{)}_{i}|s_{i}>0\}\,,$ (9)
which corresponds to the (parallelizable) objective evaluation on the obtained grid. The same formulation can support other line-search objectives that can be used instead of the loss $\mathcal{L}$, such as BER or Frame Error Rate (FER).
Training
The optimization parameters are the following: the initial $H$, the maximum number of optimization steps (if convergence is not reached), the number and quality of the data samples, the grid search length, and the number of BP iterations.
We assume that an initial $H$ is given by the user as the code to be improved. The number of optimization steps is set to 20 iterations. The training noise is sampled randomly per batch in the $\{3,\dots,7\}$ normalized SNR (i.e. $E_{b}/N_{0}$) range but can be modified according to the noise setting of interest. The number of data samples per optimization iteration is set to 4.9M for every code as sufficient statistics, and the data samples are required to have non-zero syndrome. Because of computational constraints, the number of BP iterations during training is fixed and set to 5, while other ranges or values of interest can be used instead. For faster optimization, the grid search is restricted to the first 110 smallest step sizes. Training and experiments are performed on 8x12GB GeForce RTX 2080 Ti GPUs and require 2.96 minutes on average per optimization step.

### v2

@@ Abstract
The design of optimal linear block codes capable of being efficiently decoded is of major concern, especially for short block lengths. As near capacity-approaching codes, Low-Density Parity-Check (LDPC) codes possess several advantages over other families of codes, the most notable being its efficient decoding via Belief Propagation. While many LDPC code design methods exist, the development of efficient sparse codes that meet the constraints of modern short code lengths and accommodate new channel models remains a challenge. In this work, we propose for the first time a gradient-based data-driven approach for the design of sparse codes. We develop locally optimal codes with respect to Belief Propagation decoding via the learning of the Factor graph under channel noise simulations. This is performed via a novel complete graph tensor representation of the Belief Propagation algorithm, optimized over finite fields via backpropagation and coupled with an efficient line-search method. The proposed approach is shown to outperform the decoding performance of existing popular codes by orders of magnitude and demonstrates the power of data-driven approaches for code design.
@@ 1 Introduction
Reliable digital communication is of major importance in the modern information age and involves the design of codes that can be robustly and efficiently decoded despite noisy transmission channels. Over the last half-century, significant research has been dedicated to the study of capacity-approaching Error Correcting Codes (ECC) (Shannon, 1948). Despite the initial focus on short and medium-length linear block codes (Berlekamp, 1974), the development of long channel codes (Forney, 1966; Costello & Forney, 2007) has emerged as a viable approach to approaching channel capacity (Berrou et al., 1993; MacKay, 1999; Richardson et al., 2001; Richardson & Urbanke, 2001; Arikan, 2008; Luby et al., 2001; Kudekar et al., 2011).
While the NP-hard maximum likelihood rule defines the target decoding of a given code, developing more practical solutions generally relies on theories grounded upon asymptotic analysis over conventional communication channels. However, modern communication systems rely on the design of short and medium-block-length codes (Liva et al., 2016) and the latest communication settings provide new types of channels. This is mainly due to emergent applications in the modern wireless realm requiring the transmission of short data units, such as remote command links, Internet of Things, and messaging services (De Cola et al., 2011; Boccardi et al., 2014; Paolini et al., 2015; Durisi et al., 2016; ESTI, 2021). These challenges call for the formulation of data-driven solutions, capable of adapting to various settings of interest and constraints, generally uncharted by existing theories.
The vast majority of existing machine-learning solutions to the ECC problem concentrate on the design of neural decoders. The first neural models focused on the implementation of parameterized versions of the legacy Belief Propagation (BP) decoder (Nachmani et al., 2016; 2018; Lugosch & Gross, 2017; Nachmani & Wolf, 2019; Buchberger et al., 2020). Recently, state-of-the-art learning-based de novo decoders have been introduced, borrowing from well-proven architectures from other domains. A Transformer-based decoder that incorporates the code into the architecture has been recently proposed by (Choukroun & Wolf, 2022a), outperforming existing methods by sizable margins and at a fraction of their time complexity. This architecture has been subsequently integrated into a denoising diffusion models paradigm, further improving results (Choukroun & Wolf, 2022b). Subsequently, a universal neural decoder has been proposed in (Choukroun & Wolf, 2024c), capable of unified decoding of codes from different families, lengths, and rates. Most recently and related to our work, (Choukroun & Wolf, 2024a) developed an end-to-end learning framework capable of co-learning binary linear block codes along with the neural decoder.
@@
In this work, given the ubiquity and advantages of the Belief Propagation (BP) algorithm (Pearl, 1988; Richardson et al., 2001) for sparse codes, we consider the optimization of codes with respect to BP via the learning of the underlying factor/Tanner graph. From a graphical probabilistic model perspective Koller & Friedman (2009), BP being a marginalization algorithm, a gradient-based of a score metric method is given for the structure learning of BP’s underlying Bayesian network in an end-to-end fashion. As far as we can ascertain, this is the first time a gradient-based data-driven solution is given for the design of the codes themselves based on a classical decoder. Such a solution induces a very low overhead (if any) for integration into the existing decoding solutions.
@@ 2 Related Works
Neural decoder or data-driven contributions generally focus on short and moderate-length codes for two main reasons. First, classical decoders reach the capacity of the channel for large codes, and second, the emergence of applications driven by the Internet of Things requires effective decoders for short to moderate-length codes. For example, 5G Polar codes have code lengths of 32 to 1024 (Liva et al., 2016; ESTI, 2021).
@@
Previous work on neural decoders is generally divided into two main classes: model-free and model-based. Model-free decoders employ general types of neural network architectures (Cammerer et al., 2017; Gruber et al., 2017; Kim et al., 2018b; Bennatan et al., 2018; Jiang et al., 2019a; Choukroun & Wolf, 2022a; b; 2024c; 2024b). Model-based decoders implement parameterized versions of classical Belief Propagation (BP) decoders, where the Tanner graph is unfolded into an NN in which scalar weights are assigned to each variable edge. This results in an improvement in comparison to the baseline BP method for short codes (Nachmani et al., 2016; Nachmani & Wolf, 2019; Raviv et al., 2020; 2023; Kwak et al., 2023). While model-based decoders benefit from a strong theoretical background, the architecture is overly restrictive, which generally enforces its coupling with high-complexity NN (Nachmani & Wolf, 2021). Also, the improvement gain generally vanishes for more iterations and longer codewords (Hoydis et al., 2022) and the integration cost remains very high due to both computational and memory requirements.
@@
While neural decoders show improved performance in various communication settings, there has been very limited success in the design of novel neural coding methods, which remain impracticable for deployment (O’Shea & Hoydis, 2017; Kim et al., 2018a; Jiang et al., 2019b). Recently, (Choukroun & Wolf, 2024a) provided a new differentiable way of designing binary linear block codes (i.e., parity-check matrices) for a given neural decoder also showing improved performance with classical decoders.
@@
Belief-propagation decoding has multiple advantages for LDPC codes (Gallager, 1962; Richardson & Urbanke, 2001; Richardson et al., 2001). A large number of LDPC code (parity check matrix) design techniques exist in the literature, depending on the design criterion. Among them, Gallager (Gallager, 1962) developed the first regular LDPC codes as the concatenations of permuted sub-matrices. MacKay (MacKay & Neal, 1995) demonstrated the ability of sparse codes to reach near-capacity limits via semi-randomly generated matrices. Irregular LDPC codes have been developed by (Richardson et al., 2001; Luby et al., 2001; Chung et al., 2001) where the decoding threshold can be optimized via density-evolution. Progressive Edge Growth (Hu et al., 2001; 2005) has been proposed to design large girth codes. Certain classes of LDPC array codes have been presented in (Eleftheriou & Olcer, 2002) and LDPC codes with combinatorial design constraints have been developed in (Vasic & Milenkovic, 2004). Finite geometry codes have been developed in (Lucas et al., 2000; Kou et al., 2001) and repeat-accumulate codes have been proposed by (Yang et al., 2004; Jin et al., 2000; Narayanaswami, 2001). However, the classical methods are not data-driven and are difficult to adapt to the design of codes under constrained settings of interest (e.g., short codes, modern channels, structure constraints, etc.). Most related to our work are methods for structure learning Koller & Friedman (2009) for Bayesian networks such as the Chow-Liu Algorithm Chow & Liu (1968) or search-based methods Tian (2013). Related to greedy search-based methods, Elkelesh et al. (2019) recently suggested the application of classical genetic algorithms for the discovery of better IRA codes.
@@ 3 Background
We assume a standard transmission protocol using a linear block code $C$. The code is defined by a generator matrix $G\in\{0,1\}^{k\times n}$ and the parity check matrix $H\in\{0,1\}^{(n-k)\times n}$ is defined such that $GH^{T}=0$ over the order 2 Galois field $GF(2)$. The parity check matrix $H$ entails what is known as a Tanner graph (Tanner, 1981), which consists of $n$ variable nodes and $(n-k)$ check nodes. The edges of this bipartite graph correspond to the on-bits of the matrix $H$.
The input message $m\in\{0,1\}^{k}$ is encoded by $G$ to a codeword $c\in C\subset\{0,1\}^{n}$ satisfying $Hc=H(mG)=0$ and transmitted via a Binary-Input Symmetric-Output channel, e.g., an AWGN channel. Let $y$ denote the channel output represented as $y=c_{s}+\varepsilon$, where $c_{s}$ denotes the transmission modulation of $c$ (e.g., Binary Phase Shift Keying (BPSK)), and $\varepsilon$ is random noise independent of the transmitted $c$. The main goal of the decoder $f_{H}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ conditioned on the code (i.e., $H$) is to provide a soft approximation $\hat{x}=f_{H}(y)$ of the codeword.
@@
Here, we describe ECC’s classical notation of BP with, $v\in\{1,\dots,n\}$ denotes the index corresponding to the $v^{th}$ element of the channel output $y$, for the corresponding bit $c_{v}$ we wish to recover.
@@
Here, each message indexed by the edge $e=(c,v)$ on the Tanner graph and $N(v)=\{(c,v)|H(c,v)=1\}$, i.e, the set of all edges in which $v$ participates. By definition $x^{0}=0$ such that the messages are directly determined by the vector $L$ for $k=1$.
@@
Figure 1: For the Hamming(7,4) Code: (a) parity check matrix, the induced (b) Tanner graph, and (c) the corresponding unrolled Trellis graph with two iterations, with odd layers in blue and even layers in red. In (d) we present our approach for structure learning via the learned binary weighting of the edges of the complete bipartite factor graph unlike the conventional sparse representation.
@@ 4 Method
The performance of BP is strongly tied to the underlying Tanner graph induced by the code. BP and its variants are generally implemented over a fixed sparse graph, such that the only degree of freedom resides in the number of decoding iterations. While several recent contributions (Nachmani et al., 2016; Nachmani & Wolf, 2019) aim to enhance the BP algorithm by augmenting the Trellis graph with neural networks, these approaches assume and maintain fixed codes. Here, we propose optimizing the code for the BP algorithm on a decoding setting of interest. Given a trainable binary parity check matrix $H$, we wish to obtain BP-optimized codes by solving the following parameterized optimization problem
@@
Here, $G(H)$ denotes the generator matrix defined by $H$ (i.e., $GH^{T}=0$), $\phi$ denotes the modulation function such that $c_{s}=\phi(c)$, and $\mathcal{Z}$ is the channel noise distribution. $f_{H,T}$ denotes the BP decoder built upon $H$ with $T$ iterations (sampled uniformly from a given set), $\mathcal{D}$ denotes the distance metric of interest, and $\mathcal{R}$ denotes the potential hard/soft regularization of interest, e.g., sparsity or constraints on the code structure.
@@
To obtain BP codes, we propose a structure/Tanner graph learning approach, where the bipartite graph is assumed as complete with learnable binary-weighted edges. This way, the tensor reformulation of BP weighted by $H$ allows a differentiable optimization of the Tanner graph itself. The two alternating stages of BP can now be represented in a differentiable matrix form rather than its static graph formulation, where the variable layers can be rewritten as
$Q_{ij}$ $=L_{i}+\sum_{j^{\prime}\in C_{i}\setminus j}{R_{j^{\prime}i}}\equiv L_{i}+\sum_{j^{\prime}}{R_{j^{\prime}i}H_{j^{\prime}i}}-R_{ji}\,,$ (5)
@@
$=2\text{arctanh}\left(\prod_{i^{\prime}}\left({\tanh\left(\frac{Q_{i^{\prime}j}H_{ji^{\prime}}}{2}\right)+\left(1-H_{ji^{\prime}}\right)}\right)/(\tanh\left(\frac{Q_{ij}}{2}\right)\right)\,,$
where $C_{i}$ and $V_{j}$ correspond to the non-zero elements in column $i$ and row $j$ of $H$, respectively, while the ones elements in $(1-H)\in\{0,1\}^{(n-k)\times n}$ satisfy the identity element of multiplication. Potential zero denominators have not been observed but can be handled via regularization or omission. As we can observe, BP remains differentiable with respect to $H$ as a composition of differentiable functions.
@@ 7             R = 2*atanh( prod(tmp*H+(1-H),dim=1)/tmp )
The tensor reformulation solves the major challenge of graph learning (challenge (iv)). Challenges (ii) and (iii) are also eliminated in our formulation. First, since for any given $H$ the conditional independence of error probability under symmetry (Richardson & Urbanke, 2001) is satisfied for message passing algorithms, it is enough to optimize the zero codeword only, i.e., $c=Gm=0$, removing then any dependency on $G$ in the objective (challenge (ii)). As a byproduct, we obtain that the optimization problem is invariant to the choice of modulation, whether differentiable or not (challenge (iii)).
To optimize $H$ (challenge (i)) we relax the NP-hard binary programming problems to an unconstrained objective where, given a parameter matrix $\Omega\in\mathbb{R}^{(n-k)\times n}$, we have $H\coloneqq H(\Omega)=\operatorname{bin}(\Omega)$. Here $\operatorname{bin}(\cdot)$ refers to the element-wise binarization operator implemented via the shifted straight-through-estimator (STE) (Bengio et al., 2013) defined such that
@@
where $c_{s}=\phi(c)$ denotes the modulated zero codeword and $\varepsilon_{i}$ denotes the $i^{th}$ noise sample drawn from the channel noise distribution. This objective aims to provide optimal decoding on different numbers of (variable) decoding iterations $t$.
While highly non-convex, the objective is (sub)differentiable when considering the STE definition of the gradient (Bengio et al., 2013; Yin et al., 2019) and thus optimizable via classical first-order methods. Since $H$ is binary, only changes in the sign of $\Omega$ are relevant for the optimization, so most gradient descent iterations remain ineffective in reducing the objective using conventional small learning-rate regimes. Thus, given the gradient $\nabla_{\Omega}\mathcal{L}$ computed on sufficient statistics, we propose a line-search procedure capable of finding the optimal step size.
Binary Line-Search
Conventional first-order optimization methods with small learning rate regimes have two major drawbacks with binarization (Rastegari et al., 2016; Courbariaux et al., 2016). First, they are generally slow since only gradient steps modifying the sign of the binarized tensor induce a modification of the loss. Second, they have difficulties in converging to local minima because of oscillating behavior around zero.
In general, efficient line search methods (Nocedal & Wright, 2006) assume local convexity or a smooth objective (Wolf, 1978) or, alternatively, apply exhaustive search on a given interval. Since this is not our case, we propose a novel efficient grid-search approach optimized to our binary programming setting. While classical grid search methods look for the optimal step size on handcrafted predefined sample points, in our binary setting we can search only for the step sizes inducing a flip of the sign in $\Omega$, provably limiting the maximum number of relevant grid samples to $n(n-k)$. Thus, the line-search problem is now given by
$\lambda^{*}=\operatorname*{arg\,min}_{\lambda\in\mathcal{I}_{\Omega}}\mathcal{L}(\Omega-\lambda\nabla_{\Omega}\mathcal{L}),\ \ \ \ \ \ \ \mathcal{I}_{\Omega}=\{s_{i}=\frac{(\Omega)_{i}}{(\nabla_{\Omega}\mathcal{L})_{i}}|s_{i}>0\}\,,$ (9)
which corresponds to the (parallelizable) objective evaluation on the obtained grid. The same formulation can support other more practical line-search objectives instead of the cross-entropy loss $\mathcal{L}$, such as BER or Frame Error Rate (FER).